A Normal window with a perimeter of 20 units has the shape of a rectangle surmounted by a
semi-circle. If its base is x units long and the height of the rectangle is y units long, write a mathematical model for its area in terms of x.
1
Expert's answer
2020-09-23T13:55:59-0400

Radius of semicircle"=\frac{x}{2}"

Perimeter of window"=x+2y+\frac{πx}{2}=20"

"y=10-\frac{x}{4}(2+π)"

Area of window"=xy+\frac{π}{2}\times\frac{x^2}{4}"

Substituting "y" in area, we get

Area of window"=10x-\frac{x^2}{4}(2+\frac{π}{2})"